An alloy of tin and copper has 6 pounds of copper for every 2 pounds of tin. If 200 pounds of this alloy are made, how many pounds of tin are used?
a. 25
b. 50
c. 100
d. 125

Answer by stanbon(75887)

An alloy of tin and copper has 6 pounds of copper for every 2 pounds of tin.
If 200 pounds of this alloy are made, how many pounds of tin are used?
---
6:2 is the same as 6x:2x where 2x is the amount of tin
-----
Equations:
2x + 6x = 200
8x = 200
x = 25
----
Amount of tin = 2x = 2*25 = 50 lbs.
=======================================
Cheers,
Stan H.
=======================================

a. 25
b. 50
c. 100
d. 125